OpenStudy (anonymous):

can you help me solve?x/5=x/6 +1

OpenStudy (anonymous):

Cross-multiply numerators and denominators, what do you get?

OpenStudy (anonymous):

\[\frac{ x }{ 5 }=\frac{ x }{ 6 }+1\]

OpenStudy (anonymous):

Take all containing x to one side, and 1 to the other, simplify.

OpenStudy (anonymous):

for some reason i com eou

OpenStudy (anonymous):

for some reason i come out with \[\frac{ 1 }{ -11 }\]

OpenStudy (anonymous):

is that right?